OpenStudy (anonymous):

Determine the x intercepts, zeros, and factors for the polynomial f(x)= 3x^4+3x^3-18x^2

OpenStudy (anonymous):

each term has a common factor of \(3x^2\) so the first step is to factor it out and write \[f(x)=3x^2(x^2+x-6)\]

OpenStudy (anonymous):

next step is to factor \(x^2+x-6\) do you know how to do that?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

ok, what do you get?

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

wait no 3

OpenStudy (anonymous):

i mean "what to you get when you factor?"

OpenStudy (anonymous):

more precisely, what do you get when you factor \(x^2+x-6\) ?

OpenStudy (anonymous):

I get the x intercepts

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

or the zero's

OpenStudy (anonymous):

you get the factors that you set equal to zero to find the intercepts, yes

OpenStudy (anonymous):

but my question is "how did you factor \(x^2+x-6\) ? " what are the factors?

OpenStudy (anonymous):

positive 6 ? wait okay, I add 6 to 0 and then what do I do with the x^2

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

you want to factor \(x^2+x-6\) so you have to think of two numbers whose product is \(-6\) and whose sum is 1 i.e. to numbers that, when you multiply them together you get \(-6\) and when you add them you get \(1) since -6 is negative, one of them will be positive and the other will be negative

OpenStudy (anonymous):

-3 2

OpenStudy (anonymous):

those are the zeros, yes, but actually it factors as \[x^2+x-6=(x+3)(x-2)\]

OpenStudy (anonymous):

because you switch em!

OpenStudy (anonymous):

the plus and minus

OpenStudy (anonymous):

so if \(x+3=0\) then \(x=-3\) and if \(x-2=0\) then \(x=2\)

OpenStudy (anonymous):

okay! and that's the zero"s?

OpenStudy (anonymous):

so you have three zeros

OpenStudy (anonymous):

they are \(x=2, x=-3\) and also don't forget \(x=0\)

OpenStudy (anonymous):

x=-3 x=2 and x=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay what about the x intercepts

OpenStudy (anonymous):

those are the x intercepts

OpenStudy (anonymous):

x intercept is a synonym for "zeros"

OpenStudy (anonymous):

they are the same thing??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are the factors of the polynomial the same thing as well?

OpenStudy (anonymous):

if \(f(2)=0\) that means \((0,2)\) is on the graph, i.e. that is where the graph crosses the \(x\) axis

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

\[f(x)= 3x^4+3x^3-18x^2=3x^2(x+3)(x-2)\] the "factors" are \[3x^2, x+3, x-2\]

OpenStudy (anonymous):

the "zeros" are \(\{0,-3,2\}\)

OpenStudy (anonymous):

the "x intercepts" are \((0,0), (-3,0), (2,0)\)

OpenStudy (anonymous):

oh okay, I was thinking the x intercepts where something different but I understand now

OpenStudy (anonymous):

but clearly they are related

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

yw